OpenStudy (colebenton10):

find the value of f(-9) and g(-2) if f(x)=-5x-2 and g(x)=3x^2-21x

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Replace the variable in the defining equation of the function by the given number and then simplify.

OpenStudy (skullpatrol):

$$\Huge f(-9)=-5(-9)-2$$

OpenStudy (skullpatrol):

Simplify that^ and try to do the second one on your own :-)

OpenStudy (mathstudent55):

f(x) = -5x - 2 means there is a function in x called f which is defined as y = -5x - 2. The expression f(-9) means find the y value of function f when x = -9. All you do is replace x with -9 in function f and evaluate the expression. Then you are told about a function called g, and you are told what the function is. Similarly, to find g(-2), you replace x with -2 in function g, and you evaluate the expression.

OpenStudy (mecca99):

f(x) = -5x - 2 -> f(-9)=-5x-2 \[g(x)=3x^2-21x\] -> \[g(-2)=3x^2-21x\] the new equations are f(-9)=-5x-2 and \[g(-2)=3x^2-21x\]. all you have to do is plug -9 in for the x's in the 1st equation and -2 in for the x's of the 2nd equation f(-9)=-5x-2 f(-9)=-5(-9)-2 f(-9)=45-2 f(-9)=43 \[g(-2)=3x^2-21x\] \[g(-2)=3(-2)^2-21(-2)\] g(-2)=12+42 g(-2)=54